Hi! I came across this problem in a CLEP practice test, and I'd never had any experience with it before. I'm not exactly sure how to look for tutorials because I don't know what this particular kind of problem is called. Here it is: If ln(√x) = 5 and ln(y3) = 3, what is ln(x2y)? Can you help me learn how to solve this? Thank you!

Answer by rothauserc(4718)

ln(x^(1/2)) = 5
:
eliminate ln on the left side of = by taking exp of both sides
:
x^(1/2) = e^5
:
raise both sides of = to power of two
:
x = e^10
:
ln(y^3) = 3
:
take exp of both sides of =
:
y^3 = e^3
:
y = e
:
do you mean ln(x^2y) or ln(x2y)? - I assume
:
ln(x^2y) = ln( (e^10)^2 * e) = ln(e^21) = 21
: